# An introduction to reinforcement learning for neuroscience

### v2

@@ Abstract
Reinforcement learning has a rich history in neuroscience, from early work on dopamine as a reward prediction error signal for temporal difference learning (Schultz et al., , 1997) to recent work suggesting that dopamine could implement a form of ‘distributional reinforcement learning’ popularized in deep learning (Dabney et al., , 2020). Throughout this literature, there has been a tight link between theoretical advances in reinforcement learning and neuroscientific experiments and findings. As a result, the theories describing our experimental data have become increasingly complex and difficult to navigate. In this review, we cover the basic theory underlying classical work in reinforcement learning and build up to an introductory overview of methods in modern deep reinforcement learning that have found applications in systems neuroscience. We start with an overview of the reinforcement learning problem and classical temporal difference algorithms, followed by a discussion of ‘model-free’ and ‘model-based’ reinforcement learning together with methods such as DYNA and successor representations that fall in between these two extremes. Throughout these sections, we highlight the close parallels between such machine learning methods and related work in both experimental and theoretical neuroscience. We then provide an introduction to deep reinforcement learning with examples of how these methods have been used to model different learning phenomena in systems neuroscience, such as meta-reinforcement learning (Wang et al., , 2018) and distributional reinforcement learning (Dabney et al., , 2020). Code that implements the methods discussed in this work and generates the figures is also provided.
@@ 1 Introduction
Humans and other animals learn from their experiences. Sometimes, this takes the form of explicit demonstration, as is often the case in our educational system. However, we also often have to learn from trial and error together with feedback received from the world around us – sometimes implicit and sometimes explicit. This is well illustrated by the classical case study of Pavlov’s dogs, who learned to associate a so-called ‘conditioned stimulus’ (CS; e.g. the ringing of a bell) with the availability of food shortly after (the ‘unconditioned stimulus’; US). Following a brief period of learning, the dogs would start to salivate in response to the CS in advance of any food actually being served. This suggests that the dogs had learned to associate the CS with the availability of ‘reward’ in the form of food, and that they produced an appropriate physiological response to take advantage of this food availability. Importantly, this occurred without any explicit instruction or description of the sequence of events preceding food being served. Instead, the dogs learned from experience with their environment and the presence of a salient, rewarding stimulus.
@@
Such passive stimulus-response predictions are also called ‘Pavlovian learning’ and have been commonly used in neuroscience to study learning from external rewards (Niv, , 2009). This forms a specific instantiation of the concept of ‘reinforcement learning’, which is a general term for settings where an agent learns from reward signals in the environment rather than explicit demonstration, as is the case in ‘supervised learning’. Importantly, the past decades have shown that principles of reinforcement learning can be used to explain not just behaviour, but also neural activity in biological circuits (Niv, , 2009; Botvinick et al., , 2020). An explicit neural basis of RL was initially demonstrated in foundational work by Schultz et al., (1997), which showed that the firing rates of dopaminergic neurons in the Ventral Tegmental Area (VTA) reflect the difference between expected and actual ‘value’ when animals received a juice reward following a CS consisting of a lever-press in response to a small light turning on. This provided a potential neural substrate of the classical ‘temporal difference’ learning algorithm (Schultz et al., , 1997; Sutton, , 1988), which has since been expanded to a wealth of evidence for reinforcement learning in neural dynamics (Niv, , 2009; Dabney et al., , 2020; Watabe-Uchida et al., , 2017). However, these classical algorithms are generally restricted to simple problem settings, while humans and other animals are capable of solving complex high-dimensional problems involving extended planning and complex motor control. The field of ‘deep reinforcement learning’ has recently emerged to tackle such problems in a machine learning setting, which has led to impressive results across a range of tasks (Mnih et al., , 2013; Schrittwieser et al., , 2020; Wurman et al., , 2022; Vinyals et al., , 2019). Intriguingly recent research has demonstrated that these deep RL algorithms also have parallels in both behaviour and neural dynamics (Botvinick et al., , 2020; Wang et al., , 2018; Dabney et al., , 2020; Jensen et al., , 2023; Aldarondo et al., , 2024), suggesting that neuroscience can continue to learn from advances in reinforcement learning.
@@
In this review, we provide an overview of the reinforcement learning problem and popular algorithms, with a particular focus on parallels and uses of these algorithms in neuroscience. This overview starts from classical tabular TD learning and Q-learning algorithms, which have guided neuroscientific research for decades. We then consider the important distinction between model-based and model-free reinforcement learning, as well as methods that fall somewhere in the gray area between these extremes, and discuss their neural correlates. Finally, we generalize the tabular methods to the non-linear function approximation setting and the resulting deep RL methods, which have revolutionized machine learning in recent years. We do this with a focus on methods that have had a strong influence on neuroscience to give the reader a better idea of the mathematical and computational background of recent neuroscientific findings. These include the ‘meta-reinforcement learning’ model of PFC by Wang et al., (2018) and the ‘distributional reinforcement learning’ model of VTA dopaminergic neurons by Dabney et al., (2020) in particular. We hope this review will be useful both for those who are interested in the theory underlying reinforcement learning in neuroscience and for those who want an overview of how the neuroscientific literature builds on principles from reinforcement learning theory. Throughout the paper, the focus will be on an intuitive understanding of the relevant RL methods, and explicit derivations are included only where we consider them conducive to such understanding. We refer to Sutton and Barto, (2018) for a more in-depth treatment of the underlying theory.
@@ 2 Problem setting
Here we provide a short introduction to the reinforcement learning problem in a discrete state and action space with a finite time horizon – a common setting for neuroscience experiments consisting of repeated trials or episodes in a controlled environment. In the discrete problem setting, the environment consists of states $s\in\mathcal{S}$, and the agent can take actions $a\in\mathcal{A}$. The environment is characterized by transition and reward probabilities $p(s_{t+1},r_{t}|s_{t},a_{t})$, where $r_{t}$ is the reward at time $t$. We will use $r(s,a)$ to denote either the reward when it depends deterministically on the state and action, or its expectation otherwise. We will further make the Markov assumption that the next state and reward only depend on the current state and action, $p(s_{t+1},r_{t}|s_{t},a_{t},s_{t-1},a_{t-1},...,s_{0},a_{0})=p(s_{t+1},r_{t}|s_{t},a_{t})$.
@@
Figure 1: The reinforcement learning problem and cliffworld environment. (A) An agent (here the bird) interacts with the world to maximize reward. This involves a balance between exploring potentially interesting new states (e.g. searching for food in a new field) while also exploiting states known to yield high reward (e.g. the field that had many worms yesterday). At a given point in time, the bird is in some state $s_{t}$ from which it can take an action $a_{t}$, with the probability of different actions determined by the ‘policy’ $\pi(a|s_{t})$, which is controlled by the agent. $a_{t}$ then leads to a change in the environment according to the non-controllable environment dynamics $s_{t+1},r_{t}\sim p(s_{t+1},r_{t}|s_{t},a_{t})$. Here, $r_{t}$ is the empirical ‘reward’ received by the agent, and its objective is to collect as much cumulative reward as possible. Often, reinforcement learning problems are divided into ‘episodes’, with the agent learning over the course of multiple repeated exposures to the environment. This could for example consist of the bird learning over multiple days which fields are likely to be rich in food, while minimizing the distance travelled and exposure to predators. (B) The ‘cliffworld’ environment, which will be used to demonstrate the performance and behaviour of a range of reinforcement learning algorithms in this work. The agent starts in the lower left corner (location [0, 0]), and the episode finishes when it encounters either the ‘cliff’ (dark blue) or the goal (yellow; location [9,0]). If the agent walks off the cliff, it receives a reward of -100. If it finds the goal, it receives a reward of +50. In any other state, it receives a reward of -1. Such negative rewards for ‘neutral’ actions are commonly used to encourage the agent to achieve its goal as fast as possible. The arrows indicate the ‘optimal’ policy, which takes the agent to the goal via the shortest possible route that avoids the cliff.
We can now define a trajectory $\tau=\{s_{t},a_{t},r_{t}\}_{t=0}^{T}$, where
$p(\tau)=p(s_{0})\prod_{t=0}^{T}p(s_{t+1},r_{t}|s_{t},a_{t})p(a_{t}|s_{t}).$ (1)
@@
$p(a_{t}|s_{t})$ is the probability of taking action $a_{t}$ in state $s_{t}$, which is usually controlled by the agent and denoted a policy $\pi(a_{t}|s_{t})$ (Figure 1A). The objective of the agent is to maximize the expected total discounted reward
@@
where $R_{\tau}:=\sum_{t=0}^{T}\gamma^{t}r_{t}|\tau$ and we have written $J(\pi)$ since the policy uniquely specifies $J$ in a stationary environment. In Equation 2, $\gamma$ is a ‘discount factor’, which stipulates that we should care more about immediate rewards than rewards far in the future. We can provide three interpretations for this discount factor. One is that agents intrinsically care more about immediate reward than distant reward. A second is that there is a fixed non-zero probability $(1-\gamma)$ of the current ‘episode’ or environment terminating or changing at each timestep, in which case we should weight putative future reward by the probability that we are still engaged in the task at that time. The third view is that $\gamma$ simply provides a tool for reducing the variance of our learning methods, especially in temporally extended tasks. This third view is most compatible with the fact that evaluation of RL agents after training is generally done without discounting.
Since $J(\pi$) depends on the policy of the agent, it is possible to search in the space of policies for one that maximizes $J$, which is the topic of reinforcement learning. It is often assumed that the experience $\{\tau\}$ is generated by the agent acting according to its policy, and the resulting experience is then used to update the policy in a way that increases $J(\pi)$. However, ‘off-policy’ and ‘offline’ reinforcement learning methods also exist, where the agent learns on the basis of experience generated by a policy different from $\pi$ (Levine et al., , 2020; Section 11.2). This is important for biological organisms, where learning can happen from observing other individuals or ‘offline’ during sleep and rest.
@@ 3 Temporal difference learning
A simple way to mazimize reward in an environment is to learn the ‘value’ of different states, and then move towards states with high value. The potential importance of such an algorithm for neuroscience is evident from the value-seeking behaviour of many organisms, and the widespread findings of neural ‘codes’ for value across the brain (Schultz et al., , 1992; Padoa-Schioppa and Assad, , 2006; Rushworth et al., , 2011). This leaves the question of how such value codes can be learned in a biologically plausible setting.
A simple answer to this question takes the form of the classical ‘temporal difference learning’ algorithm (Sutton, , 1988; Sutton and Barto, , 2018). This involves defining a value function for a given state $s$ and policy $\pi$, which quantifies the expected future reward when following $\pi$ starting from $s$:
@@
$V^{\pi}(s)$ $=r(s)+\sum_{s^{\prime}}p_{\pi}(s_{t+1}=s^{\prime}|s_{t}=s)\mathbb{E}_{\tau\sim p_{\pi}(\tau)}\left[\sum_{t^{\prime}=t+1}\gamma^{t^{\prime}-t}r_{t^{\prime}}|s_{t+1}=s^{\prime}\right]$ (4)
$=r(s)+\gamma\sum_{s^{\prime}}p_{\pi}(s_{t+1}=s^{\prime}|s_{t}=s)V^{\pi}(s^{\prime}),$ (5)
where $p_{\pi}(s_{t+1}=s^{\prime}|s_{t}=s)=\sum_{a}\pi(a|s)p(s_{t+1}=s^{\prime}|s_{t}=s,a_{t}=a)$ is the probability of transitioning from $s$ to $s^{\prime}$ under $\pi$. Importantly, Equation 5 would not hold if $V^{\pi}(s)$ was not the true value function (Sutton and Barto, , 2018). When learning an approximate value function $V(s)$, we can therefore use this bootstrapped self-consistency relation as an objective function:
$\mathcal{L}\propto\left(V(s)-\left(r(s)+\gamma\mathbb{E}_{p_{\pi}(s^{\prime}|s)}\left[V(s^{\prime})\right]\right)\right)^{2},$ (6)
@@
$\propto-V(s)+r(s)+\gamma\mathbb{E}_{p_{\pi}(s^{\prime}|s)}\left[V(s^{\prime})\right]$ (8)
@@
The last line approximates the expected update with a single sample corresponding to the states and reward actually experienced. As more experience is collected and many small gradient steps are taken according Equation 9, these single-sample estimates average out to the expectation in Equation 8 (Figure 2A). Variants of this algorithm can also learn about multiple past states at once using the notion of eligibility traces (Sutton and Barto, , 2018). However, Equation 9 is the canonical temporal difference learning rule (Sutton, , 1988), and it leads to learning dynamics where the temporal difference error $\delta_{t}:=-V(s_{t})+r_{t}+\gamma V(s_{t+1})$ gradually propagates from the rewarding state to prior states that predict the upcoming reward (Figure 2C).
@@
This gradual propagation of prediction errors from the reward to its predecessors has been of great interest in neuroscience. In particular, classical work by Schultz et al., (1997) demonstrated a similar pattern of neural activity in dopaminergic VTA neurons, which formed the foundation of a now well-established theory that dopamine provides a biological reward prediction error signal that drives learning (Niv, , 2009; Watabe-Uchida et al., , 2017). At a behavioural level, this is supported by experiments showing that artificial stimulation of dopamine neurons can be a strong driver of learning (Olds and Milner, , 1954; Tsai et al., , 2009; Steinberg et al., , 2013). The simple narrative of dopamine as a reward prediction error has also been challenged in recent years (Coddington and Dudman, , 2019; Howe et al., , 2013; Horvitz, , 2000), which has led to theories of dopamine as a more general prediction error for both value but also other quantities like salience and movement (Kakade and Dayan, , 2002; Gershman et al., , 2024).
@@
Figure 2: Temporal difference learning. (A) Value functions aquired through temporal difference learning (Equation 9) while acting according to either a random (top) or an optimal (bottom) policy. These simulations were performed with a random start state in the cliffworld environment to ensure full coverage of the space. Dark blue indicates negative expected reward (-100) and yellow indicates positive expected reward (+50). These simulations used a learning rate of $\alpha=0.05$ and no temporal discounting ($\gamma=1$). Under the random policy, states near the cliff have low value even if they are close to the goal, since the agent often falls off the cliff from there. Under the optimal policy, all states have high expected reward, since the agent always reaches the goal. States nearer the goal have higher value than those further away, although this is hard to distinguish on this color scale. (B) Empirical reward as a function of episode number for a TD-learning agent that acts according to Equation 11 while updating its value estimates according to Equation 9. For this agent, action selection assumes access to a ‘one-step’ world model in order to evaluate the consequence of each putative action. The agent gradually converges to an optimal policy. Parameters for the agent are as in (A), except that the start state is always the lower left corner. (C) TD error (Equation 9) as a function of the step number along the optimal path for the agent in (B) at different stages of learning (green to blue). This TD signal gradually propagates backwards from the reward to preceding states, mirroring biological recordings of dopamine activity (Schultz et al., , 1997). (D) Value function learned by a greedy TD agent as in (B), plotted either early (top) or late (bottom) in training. Early in training, the agent has learned that the cliff is bad but doesn’t know where the goal is or how to get there. Late in training, the agent has learned a value function that locally resembles the optimal value function from (A), while it has not learned the value of distant states that are rarely or never visited from the start state. This is a potential shortcoming of ‘greedy’ agents that can easily converge to a sub-optimal local maximum in more complicated environments. For this analysis, we used a high learning rate of $\alpha=0.5$ to make the early TD updates larger and therefore more visible.
@@
Much classical work in neuroscience has focused on value learning in Pavlovian conditioning tasks, but animals in natural environments also have to take actions on the basis of this information. However, having learned a value function, it is simple to use it for action selection if we can estimate $r(s,a)$ and $p(s^{\prime}|s,a)$. In this case, the expected reward associated with action $a$ can be written as
$\mathbb{E}\left[\sum_{t^{\prime}=t}\gamma^{t^{\prime}-t}r_{t^{\prime}}|s_{t}=s,a_{t}=a\right]=r(s,a)+\gamma\sum_{s^{\prime}}p(s^{\prime}|s,a)V(s^{\prime}).$ (10)
We can then perform action selection by choosing the action with the highest expected reward,
$a^{*}(s)=\text{argmax}_{a}\left(r(s,a)+\gamma\sum_{s^{\prime}}p(s^{\prime}|s,a)V(s^{\prime})\right).$ (11)

### v3

@@ Abstract
Reinforcement learning (RL) has a rich history in neuroscience, from early work on dopamine as a reward prediction error signal [97] to recent work proposing that the brain could implement a form of ‘distributional reinforcement learning’ popularized in machine learning [20]. There has been a close link between theoretical advances in reinforcement learning and neuroscience experiments throughout this literature, and the theories describing the experimental data have therefore become increasingly complex. Here, we provide an introduction and mathematical background to many of the methods that have been used in systems neroscience. We start with an overview of the RL problem and classical temporal difference algorithms, followed by a discussion of ‘model-free’, ‘model-based’, and intermediate RL algorithms. We then introduce deep reinforcement learning and discuss how this framework has led to new insights in neuroscience. This includes a particular focus on meta-reinforcement learning [123] and distributional RL [20]. Finally, we discuss potential shortcomings of the RL formalism for neuroscience and highlight open questions in the field. Code that implements the methods discussed and generates the figures is also provided.
@@ 1 Introduction
Humans and other animals learn from their experiences. Sometimes, this takes the form of explicit demonstration, as is often the case during formal education. However, we also often have to learn from trial and error together with feedback received from the world around us – sometimes implicit and sometimes explicit. This is well illustrated by the classical case study of Pavlov’s dogs, who learned to associate a so-called ‘conditioned stimulus’ (CS; e.g. the ringing of a bell) with the availability of food shortly after (the ‘unconditioned stimulus’; US). Following a brief period of learning, the dogs would start to salivate in response to the CS in advance of any food actually being served. This suggests that the dogs had learned to associate the CS with the availability of ‘reward’ in the form of food, and that they produced an appropriate physiological response to take advantage of this food availability. Importantly, this occurred without any explicit instruction or description of the sequence of events preceding food being served. Instead, the dogs learned from experience with their environment and the presence of a salient, rewarding stimulus.
@@
Such passive stimulus-response predictions are also called ‘Pavlovian learning’ and have been commonly used in neuroscience to study learning from external rewards [76]. This forms a specific instantiation of the concept of ‘reinforcement learning’, which is a general term for settings where an agent learns from reward signals in the environment rather than explicit demonstration, as is the case in ‘supervised learning’. Importantly, the past decades have shown that principles of reinforcement learning can be used to explain not just behaviour, but also neural activity in biological circuits [76, 14]. An explicit neural basis of RL was initially demonstrated in foundational work by Schultz et al. [97], which showed that the firing rates of dopaminergic neurons in the Ventral Tegmental Area (VTA) reflected the difference between expected and actual ‘value’ when animals received a juice reward following a CS consisting of a lever-press in response to a small light turning on. This provided a potential neural substrate of the classical ‘temporal difference’ learning algorithm [97, 107], which has since been expanded to a wealth of evidence for reinforcement learning in neural dynamics [76, 20, 125]. However, these classical algorithms are generally restricted to simple problem settings, while humans and other animals are capable of solving complex high-dimensional problems involving extended planning and fine motor control. The field of ‘deep reinforcement learning’ has recently emerged to tackle such problems in a machine learning setting, which has led to impressive results across a range of tasks [71, 95, 131, 121]. Intriguingly, recent research has demonstrated that these deep RL algorithms also have parallels in both behaviour and neural dynamics [14, 123, 20, 47, 4], suggesting that neuroscience can continue to learn from advances in reinforcement learning.
@@
In this review, we provide an overview of the reinforcement learning problem and popular algorithms, with a particular focus on parallels and uses of these algorithms in neuroscience. This overview starts from classical tabular TD learning and Q-learning algorithms, which have guided neuroscience research for decades. We then consider the important distinction between model-based and model-free reinforcement learning, as well as methods that fall somewhere in the gray area between these extremes, and discuss their neural correlates. Finally, we generalize the tabular methods to the non-linear function approximation setting and the resulting deep RL methods, which have revolutionized machine learning in recent years. We do this with a focus on methods that have had a strong influence on neuroscience to give the reader a better idea of the mathematical and computational background of recent neuroscientific findings. These include the ‘meta-reinforcement learning’ model of PFC by Wang et al. [123] and the ‘distributional reinforcement learning’ model of VTA dopaminergic neurons by Dabney et al. [20] in particular. We hope this review will be useful both for those who are interested in the theory underlying reinforcement learning in neuroscience and for those who want an overview of how the neuroscience literature builds on principles from reinforcement learning theory. Throughout the paper, the focus will be on an intuitive understanding of the relevant RL methods, and explicit derivations are included only where we consider them conducive to such understanding. We refer to Sutton and Barto [108] for a more in-depth treatment of the underlying theory.
@@ 2 Problem setting
Here we provide a short introduction to the reinforcement learning problem in a discrete state and action space with a finite time horizon – a common setting for neuroscience experiments consisting of repeated trials or episodes in a controlled environment. In this setting, the environment consists of states from a discrete set $s\in\mathcal{S}={\{s_{1},s_{2},\ldots\}}_{1}^{|\mathcal{S}|}$, and the agent can take actions $a\in\mathcal{A}={\{a_{1},a_{2},\ldots\}}_{1}^{|\mathcal{A}|}$. The environment is characterized by transition and reward probabilities $p(s_{t+1},r_{t}|s_{t},a_{t})$, where $r_{t}$ is the reward at time $t$. We will use $r(s,a)$ to denote either the reward when it depends deterministically on the state and action, or its expectation otherwise. We will further make the Markov assumption that the next state and reward only depend on the current state and action, $p(s_{t+1},r_{t}|s_{t},a_{t},s_{t-1},a_{t-1},\ldots,s_{0},a_{0})=p(s_{t+1},r_{t}|s_{t},a_{t})$.
@@
Figure 1: The reinforcement learning problem and cliffworld environment. (A) An agent (here the bird) interacts with the world to maximize reward. This involves a balance between exploring potentially interesting new states (e.g. searching for food in a new field) while also exploiting states known to yield high reward (e.g. the field that had many worms yesterday). At a given point in time, the bird is in some state $s_{t}$ from which it can take an action $a_{t}$, with the probability of different actions determined by the ‘policy’ $\pi(a|s_{t})$, which is controlled by the agent. $a_{t}$ then leads to a change in the environment according to the non-controllable environment dynamics $s_{t+1},r_{t}\sim p(s,r|s_{t},a_{t})$. Here, $r_{t}$ is the empirical ‘reward’ received by the agent, and its objective is to collect as much cumulative reward as possible. Often, reinforcement learning problems are divided into ‘episodes’, with the agent learning over the course of multiple repeated exposures to the environment. This could for example consist of the bird learning over multiple days which fields are likely to be rich in food, while minimizing the distance travelled and exposure to predators. (B) The ‘cliffworld’ environment, which will be used to demonstrate the performance and behaviour of a range of reinforcement learning algorithms in this work. The agent starts in the lower left corner (location [0, 0]), and the episode finishes when it encounters either the ‘cliff’ (dark blue) or the goal (yellow; location [9,0]). If the agent walks off the cliff, it receives a reward of -100. If it finds the goal, it receives a reward of +50. In any other state, it receives a reward of -1. Such negative rewards for ‘neutral’ actions are commonly used to encourage the agent to achieve its goal as fast as possible. The arrows indicate the ‘optimal’ policy, which takes the agent to the goal via the shortest possible route that avoids the cliff.
We can now define a trajectory $\tau={\{s_{t},a_{t},r_{t}\}}_{t=0}^{T}$. The probability of a trajectory occuring is
$p_{\pi}(\tau)=p(s_{0})\prod_{t=0}^{T}p(s_{t+1},r_{t}|s_{t},a_{t})\pi(a_{t}|s_{t}).$ (1)
@@
$p_{\pi}(\tau)$ depends on the policy of the agent, $\pi(a|s)$, which specifies the probability of taking action $a$ in state $s$ (Figure 1A). The objective is to learn a policy that maximizes the expected total discounted reward
@@
where $R_{\tau}:=\sum_{t=0}^{T}\gamma^{t}r_{t}|\tau$, and we have written $J(\pi)$ since the policy uniquely specifies $J$ in a stationary environment. In Equation 2, $\gamma$ is a ‘discount factor’, which stipulates that we should care more about immediate rewards than rewards far in the future. We can provide three interpretations for this discount factor. One is that agents intrinsically care more about immediate reward than distant reward. A second is that there is a fixed non-zero probability $(1-\gamma)$ of the current ‘episode’ or environment terminating or changing at each timestep, in which case we should weight putative future reward by the probability that we are still engaged in the task at that time. The third view is that $\gamma$ simply provides a tool for reducing the variance of our learning methods, especially in temporally extended tasks. This third view is most compatible with the fact that evaluation of RL agents after training is generally done without discounting.
Since $J(\pi$) depends on the policy of the agent, it is possible to search in the space of policies for one that maximizes $J$, which is the topic of reinforcement learning. It is often assumed that the experience $\{\tau\}$ is generated by the agent acting according to its policy, and the resulting experience is then used to update the policy in a way that increases $J(\pi)$. However, ‘off-policy’ and ‘offline’ reinforcement learning methods also exist, where the agent learns on the basis of experience generated by a policy different from $\pi$ (57; Section 11.2). This can be either an ‘old’ version of the agent itself, when it acted according to a different policy, or data generated by an entirely different agent. Off-policy learning is important for biological organisms, where learning can happen ‘offline’ during sleep after initial data collection during wake, or from observing other individuals (also the topic of ‘imitation learning’; Section 11.3).
@@ 3 Temporal difference learning
A simple way to maximize reward in an environment is to learn the ‘value’ of different states, and then move towards states with high value. The potential importance of such an algorithm for neuroscience is evident from the value-seeking behaviour of many organisms, and the widespread findings of neural ‘codes’ for value across the brain [96, 83, 93]. This leaves the question of how such value codes can be learned in a biologically plausible setting.
One answer to this question takes the form of the classical ‘temporal difference learning’ algorithm [107, 108]. This involves defining a value function for a given state $s$ and policy $\pi$, which quantifies the expected future reward when following $\pi$ starting from $s$:
@@
$V^{\pi}(s)$ $=r_{\pi}(s)+\sum_{s^{\prime}}p_{\pi}(s_{t+1}=s^{\prime}|s_{t}=s)\mathbb{E}_{\tau\sim p_{\pi}(\tau)}\left[\sum_{t^{\prime}=t+1}\gamma^{t^{\prime}-t}r_{t^{\prime}}|s_{t+1}=s^{\prime}\right]$ (4)
$=r_{\pi}(s)+\gamma\sum_{s^{\prime}}p_{\pi}(s_{t+1}=s^{\prime}|s_{t}=s)V^{\pi}(s^{\prime}),$ (5)
where $p_{\pi}(s_{t+1}=s^{\prime}|s_{t}=s)=\sum_{a}\pi(a|s)p(s_{t+1}=s^{\prime}|s_{t}=s,a_{t}=a)$ is the probability of transitioning from $s$ to $s^{\prime}$ under $\pi$, and $r_{\pi}(s)=\mathbb{E}_{a\sim\pi}[r(s,a)]$ is the expected reward in state $s$, averaged over actions. Importantly, Equation 5 would not hold if $V^{\pi}(s)$ was not the true value function [108]. When learning an approximate value function $V(s)$, we can therefore use this bootstrapped self-consistency relation as an objective function:
$\mathcal{L}\propto{\left(V(s)-\left(r_{\pi}(s)+\gamma\mathbb{E}_{p_{\pi}(s^{\prime}|s)}\left[V(s^{\prime})\right]\right)\right)}^{2},$ (6)
@@
$\propto-V(s)+r_{\pi}(s)+\gamma\mathbb{E}_{p_{\pi}(s^{\prime}|s)}\left[V(s^{\prime})\right]$ (8)
@@
The last line approximates the expected update with a single sample corresponding to the states and reward actually experienced. As more experience is collected and many small gradient steps are taken according Equation 9, these single-sample estimates average out to the expectation in Equation 8 (Figure 2A). Variants of this algorithm can also learn about multiple past states at once using the notion of eligibility traces [108]. However, Equation 9 is the canonical temporal difference learning rule [107], and it leads to learning dynamics where the temporal difference error $\delta_{t}:=-V(s_{t})+r_{t}+\gamma V(s_{t+1})$ gradually propagates from the rewarding state to prior states that predict the upcoming reward (Figure 2C).
@@
This gradual propagation of prediction errors from the reward state to its predecessors has been of great interest in neuroscience. In particular, classical work by Schultz et al. [97] demonstrated a similar pattern of neural activity in dopaminergic VTA neurons, which formed the foundation of a now well-established theory that dopamine provides a biological reward prediction error signal that drives learning [76, 125]. At a behavioural level, this is supported by experiments showing that artificial stimulation of dopamine neurons can be a strong driver of learning [80, 117, 104]. The simple narrative of dopamine as a reward prediction error has also been challenged in recent years [19, 44, 43], which has led to theories of dopamine as a more general prediction error for both value but also other quantities like salience and movement [50, 37].
@@
Figure 2: Temporal difference learning. (A) Value functions aquired through temporal difference learning (Equation 9) while acting according to either a random (top) or an optimal (bottom) policy. These simulations were performed with a random start state in the cliffworld environment to ensure full coverage of the space. Dark blue indicates negative expected reward (-100) and yellow indicates positive expected reward (+50). These simulations used a learning rate of $\alpha=0.05$ and no temporal discounting ($\gamma=1$). Under the random policy, states near the cliff have low value even if they are close to the goal, since the agent often falls off the cliff from there. Under the optimal policy, all states have high expected reward, since the agent always reaches the goal. States nearer the goal have slightly higher value than those further away. (B) Empirical reward as a function of episode number for a TD-learning agent that acts according to Equation 11 while updating its value estimates according to Equation 9. For this agent, action selection assumes access to a ‘one-step’ world model in order to evaluate the consequence of each putative action. The agent gradually converges to an optimal policy. Parameters for the agent are as in (A), except that the start state is always the lower left corner. (C) TD error (Equation 9) as a function of the step number along the optimal path for the agent in (B) at different stages of learning (green to blue). This TD signal gradually propagates backwards from the reward to preceding states, mirroring biological recordings of dopamine activity [97]. (D) Value function learned by a greedy TD agent as in (B), plotted either early (top) or late (bottom) in training. Early in training, the agent has learned that the cliff is bad but doesn’t know where the goal is or how to get there. Late in training, the agent has learned a value function that locally resembles the optimal value function from (A), while it has not learned the value of distant states that are rarely or never visited from the start state. This is a potential shortcoming of ‘greedy’ agents that can easily converge to a sub-optimal local maximum in more complicated environments. For this analysis, we used a high learning rate of $\alpha=0.5$ to make the early TD updates larger and therefore more visible.
@@
Much classical work in neuroscience has focused on value learning in Pavlovian conditioning tasks, but animals in natural environments also have to take actions on the basis of this information. However, after learning a value function, it can be used for optimal action selection if we can estimate $r(s,a)$ and $p(s^{\prime}|s,a)$. We can then compute a ‘state-action value’ $Q^{\pi}(s,a)$, defined as the expected discounted future reward associated with taking action $a$ in state $s$ and then following policy $\pi$:
$Q^{\pi}(s,a):=\mathbb{E}\left[\sum_{t^{\prime}=t}\gamma^{t^{\prime}-t}r_{t^{\prime}}|s_{t}=s,a_{t}=a\right]=r(s,a)+\gamma\sum_{s^{\prime}}p(s^{\prime}|s,a)V^{\pi}(s^{\prime}).$ (10)
Approximating the true value function $V^{\pi}$ by the learned value function $V$ in Equation 10 yields an approximate state-action value $Q$, which can be used to choose the action with the highest expected reward,
$a^{*}(s)=\text{argmax}_{a}Q(s,a).$ (11)